OpenStudy (anonymous):

Simplify sec^2 x - 2 - tan^2 x

OpenStudy (anonymous):

Sec^2 - (sin^2 + cos^2) - (sin^2 + cos^2) - (sin/cos)^2 ?

OpenStudy (anonymous):

( Sec - (sin + cos) - (sin + cos) - (sin/cos) )^2

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\(\large\color{slate}{ \sec^2x -2-\tan^2x }\). And you need to simplify this?

OpenStudy (anonymous):

sec^2 x - 2 = tan^2 x sorry

OpenStudy (anonymous):

Then find the solutions

OpenStudy (solomonzelman):

\(\large\color{slate}{ \sec^2x -2=\tan^2x }\)

OpenStudy (anonymous):

Which shouldn't be too hard after the problem is simplified

OpenStudy (anonymous):

Yes that is the problem

OpenStudy (solomonzelman):

apply: \(\large\color{red}{ \sec^2x=\tan^2x+1 }\)

OpenStudy (anonymous):

Ok I will

OpenStudy (anonymous):

tan^2 x - 1 - tan^2 x ?

OpenStudy (anonymous):

Then -1

OpenStudy (solomonzelman):

\(\large\color{slate}{ \tan^2x -1=\tan^2x }\) ?

OpenStudy (solomonzelman):

that is what you meant?

OpenStudy (anonymous):

tan^2 x - 1 - tan^2 x = 0

OpenStudy (solomonzelman):

yes, and simplify the left side

OpenStudy (anonymous):

How?

OpenStudy (solomonzelman):

\(\large\color{slate}{ \tan^2x -1-\tan^2x=0 }\) something cancels.....

OpenStudy (anonymous):

Tan^2

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

So -1 = 0

OpenStudy (anonymous):

Then -(sin^2 x + cos^2 x) = 0?

OpenStudy (solomonzelman):

no, once you get \(\large\color{slate}{ -1=0 }\), that means that there is no solution

OpenStudy (anonymous):

Thanks! It makes sense.

OpenStudy (solomonzelman):

Oh, I opened my profile and saw that you finished typing and this last reply. In my post-tab it said that you were typing still..... anyways, you welcome!